### v1

@@ 3.4.2 Uniform-limb transit model
When fitting the data assuming or forcing the planet to have uniform limbs, we use the transit modelling package batman [54]. This code models the planetary disk in transit as a perfect circle, where the radius of this circle varies only with wavelength. We allowed the orbital parameters to be free for our ancillary data fit described below, and fixed them to the result of that fit (Table 1) when analyzing our JWST/NIRCam F322W2 spectroscopic data.
@@ 3.4.3 Ancillary Data Fits
We ran this sampling for 10,500 steps, which was sufficient for all parameters to converge. The resulting best-fit values and derived parameter values are listed in Table 1. The data and best-fit models are plotted in Figure 7 of the Extended Data, and we highlight our constraints on WASP-107b’s orbital parameters in Figure 8 of the Extended Data.
@@ 3.4.5 WASP-107b’s Obliquity and Orbital Misalignment
From our fit of the radial velocity data, we measure WASP-107b’s eccentricity to be e = 0.05 0.01 (Table 1), consistent with that of ref.[5] which analyzed the same data. Neither this eccentricity, or any parameter in our fit, is affected by the Rossiter-McLaughlin (RM) effect. The RM effect only affects radial velocity measurements made during the planet’s transit, and all of the measurements we use were made outside of WASP-107b’s transit. Also, the measurement of the RM effect in this system by ref.[55] was done by assuming and fixing WASP-107b’s orbital parameters to those derived by ref.[5], with whom we are fully consistent. Relatedly, the detailed models of ref.[56] show that the RM effect does not bias transmission spectra observed by JWST, especially at the very low spectral resolution that we bin our data to.
@@ 3.5 Constraining WASP-107b’s time of conjunction
As described in Section 3.4.3, we fit the ancillary transit and radial velocity observations described in Sections 1.1 and 3.3 together simultaneously. The goal of this fit was to precisely constrain WASP-107b’s orbit before fitting our NIRCam/F322W2 spectroscopic data. These orbital parameters must be known precisely a priori to enable accurately separating the contribution of each limb to the observed transmission signal. Uncertainty in the time of conjunction (), in particular, is degenerate with the effect of limb asymmetry on the planet’s transit light curve [7, 21, 9, 22]. Using these ancillary data, we were able to strongly constrain each of WASP-107b’s main orbital parameters (Table 1). We highlight the resulting posterior distributions of several primary orbital parameters from this fit in Figure 8 of the Extended Data. Our best-fit values are consistent with previous measurements from various other datasets [2, 1, 4, 5, 63] and, most importantly, we achieve very high precision on WASP-107b’s . At the epoch of our NIRCam F322W2 observation, we measure WASP-107b’s to a 1- precision of 0.70 seconds. This tight constraint is enabled by the wide wavelength coverage of these ancillary data, and the unprecedented combination of high photometric precision and fast integration cadence of the included JWST data.
@@
This approach to deriving assumes that limb asymmetry varies chromatically, though this cannot be empirically known a priori. However, the fact that we derive different morning and evening spectra in one data set (Figure 2) is proof that the exact same limb asymmetry is not uniformly present at all wavelengths. For instance, when fitting our broadband NIRCam F322W2 data by itself assuming a uniform-limb model and allowing to freely vary, we get a best-fit that is 6.64 1.91 seconds later than the joint-fit value (Figure 9). If the exact same limb asymmetry were present at all wavelengths and the biases did not cancel out as previously described, then our joint-fit would have given a time offset by this same value. If we then applied this offset time in our catwoman fits to the F322W2 spectroscopic data, we would not have observed any significant limb asymmetry. In this case, the assumed time of conjunction would be biased by an amount corresponding to the underlying asymmetry, which would have erased the asymmetry’s signal.
@@
To further evaluate our joint-fit’s effectiveness, we tested fitting each ancillary data set individually using a uniform-limb model and allowing to freely vary. We then compared the difference between each individually fit and the joint-fit . As just mentioned, when fitting the broadband JWST/NIRCam F322W2 data, we get a best-fit that 6.64 1.91 seconds later than the joint-fit result from the broadband light curve, and we also see chromatic offsets in the best-fit shown in Figure 9 from the spectroscopic data. Both the SOAR and Spitzer/IRAC Ch.2 observations had differences consistent with zero (7.6 19.7 s and 4.42 7.39 s, respectively). The TESS data, which overlaps with the SOAR bandpass, had a best-fit offset of -33.5 19.0 s. This is negative and non-zero at over 1 significance, but it is not clear whether such large offset is real given the sparse 30-minute cadence of the data. The JWST/NIRCam F210M data, which was observed simultaneously with the NIRCam F322W2 data with the same cadence, had an offset of -1.11 1.06 s. This is also negative but nearly consistent with zero. Finally, the JWST/MIRI LRS data had a difference of 2.62 1.28 s which is non-zero by just over 2. Due to their superior data quality, our joint-fit is most effected by the JWST/NIRCam F210M and JWST/MIRI LRS observations. These individual fits show that both data sets may also be subject to some limb asymmetry, though not to the same degree as the JWST/NIRCam F322W2 data, and potentially of opposite polarity. Our multi-wavelength approach will therefore work as described to mitigate these individual biases. All together, these fits underscore the importance of combining data at multiple wavelengths. The mean offset of these individual fit values, weighted by their uncertainties (), is 0.41 s which is well within the joint-fit value’s uncertainty. Even if we ignore the largest offset, which was the TESS value, the mean offset is still only 0.47 s. To the best of our current knowledge, the we derive from our joint-fit is as accurate and bias-free as possible.
@@ 3.6 Forward Modeled Transmission Spectra
From the grid of models described above, we used an atmospheric metallicity of 10 times the solar value and carbon-to-oxygen ratio of 0.35 based on the best-fit retrieval values on HST/WFC3 and JWST/MIRI observations of this planet from ref.[26]. These values are also consistent with the additional retrievals in Welbanks, et al. (in review). We expect the metallicity and carbon-to-oxygen ratio to be globally uniform quantities, and do not vary them between the models for each limb. In this work, we do not seek to make specific, precise constraints on global properties, but rather develop an understanding for what may be causing the observed difference between the evening and morning limbs. To do this, we performed a grid fit over the heat redistribution parameter, over its full range, and the gray cloud opacity, ranging from 10-2.58 to 10-0.78 cm2 g-1. This heat redistribution parameter is effectively the temperature at the terminator, related by . We computed the value between the grid model and the observed morning and evening spectra individually, following the -minimization technique of refs.[40, 41].
@@
Acknowledgements
Declarations
Funding: T.P.G. and T.J.B. acknowledge funding from NASA in WBS
411672.07.05.05.03.02. M.M., E.S., and M.R. acknowledge funding from NASA Goddard Spaceflight Center via NASA contract NAS5-02105. -
Conflict of interest/Competing interests: None
Ethics approval: Not applicable
Consent to participate: Not applicable.
Consent for publication: Not applicable.
Availability of data and materials: The JWST/NIRCam and JWST/MIRI data used in this analysis have completed their one-year proprietary period and are currently available in the Mikulski Archive for Space Telescopes (MAST). The other data used in this analysis are also publicly available in MAST, or can be provided upon request.
Code availability: The three reduction pipelines used in this work (Eureka!, t-shirt, and Pegasus) are all either currently open-source or are planned to be made open-source. The codes of each are hosted on Github, and links for each may be found in the Methods section.
Authors’ contributions: M.M. led this analysis including formulating the idea, performing the data and supporting analyses, and was the primary author of the paper. T.G.B. contributed the Pegasus reduction of the NIRCam data, E.S. contributed the t-shirt reduction, and T.J.B. contributed the Eureka! reduction of the NIRCam F322W2 spectra and the broadband reduction of the MIRI/LRS data. T.J.B., E.S., and T.G.B. also contributed to the interpretation of our results and the writing of the paper. M.L. contributed the modeling analysis using ScCHIMERA. T.P.G. selected the object, devised the observing strategy, and contributed to the analysis. V.P., E.R., J.F., M.L., and L.W. each contributed to the methodology and interpretation of the results. M.R. designed the JWST/NIRCam instrument used for the observations, and provided the observing time necessary for completing these observations.

### v2

@@ 3.4.2 Uniform-limb transit model
When fitting the data assuming or forcing the planet to have uniform limbs, we use the transit modelling package batman [56]. This code models the planetary disk in transit as a perfect circle, where the radius of this circle varies only with wavelength. We allowed the orbital parameters to be free for our ancillary data fit described below, and fixed them to the result of that fit (Extended Data Table 1) when analyzing our JWST/NIRCam F322W2 spectroscopic data.
@@ 3.4.3 Ancillary Data Fits
We ran this sampling for 10,500 steps, which was sufficient for all parameters to converge. The resulting best-fit values and derived parameter values are listed in Extended Data Table 1. The data and best-fit models are plotted in Supplementary Figure 2, and we highlight our constraints on WASP-107b’s orbital parameters in Supplementary Figure 3.
@@ 3.4.5 WASP-107b’s Obliquity and Orbital Misalignment
From our fit of the radial velocity data, we measure WASP-107b’s eccentricity to be e = 0.05 0.01 (Extended Data Table 1), consistent with that of ref.[5] which analyzed the same data. Neither this eccentricity, or any parameter in our fit, is affected by the Rossiter-McLaughlin (RM) effect. The RM effect only affects radial velocity measurements made during the planet’s transit, and all of the measurements we use were made outside of WASP-107b’s transit. Also, the measurement of the RM effect in this system by ref.[57] was done by assuming and fixing WASP-107b’s orbital parameters to those derived by ref.[5], with whom we are fully consistent. Relatedly, the detailed models of ref.[58] show that the RM effect does not bias transmission spectra observed by JWST, especially at the very low spectral resolution that we bin our data to.
@@ 3.5 Constraining WASP-107b’s time of conjunction
As described in Section 3.4.3, we fit the ancillary transit and radial velocity observations described in Sections 1.1 and 3.3 together simultaneously. The goal of this fit was to precisely constrain WASP-107b’s orbit before fitting our NIRCam/F322W2 spectroscopic data. These orbital parameters must be known precisely a priori to enable accurately separating the contribution of each limb to the observed transmission signal. Uncertainty in the time of conjunction (), in particular, is degenerate with the effect of limb asymmetry on the planet’s transit light curve [7, 21, 9, 22]. Using these ancillary data, we were able to strongly constrain each of WASP-107b’s main orbital parameters (Extended Data Table 1). We highlight the resulting posterior distributions of several primary orbital parameters from this fit in Supplementary Figure 3. Our best-fit values are consistent with previous measurements from various other datasets [2, 1, 4, 5, 65] and, most importantly, we achieve very high precision on WASP-107b’s . At the epoch of our NIRCam F322W2 observation, we measure WASP-107b’s to a 1- precision of 0.70 seconds. This tight constraint is enabled by the wide wavelength coverage of these ancillary data, and the unprecedented combination of high photometric precision and fast integration cadence of the included JWST data.
@@
This approach to deriving assumes that limb asymmetry varies chromatically, though this cannot be empirically known a priori. However, the fact that we derive different morning and evening spectra in one data set (Figure 2) is proof that the exact same limb asymmetry is not uniformly present at all wavelengths. For instance, when fitting our broadband NIRCam F322W2 data by itself assuming a uniform-limb model and allowing to freely vary, we get a best-fit that is 6.64 1.91 seconds later than the joint-fit value (Supplementary Figure 4). If the exact same limb asymmetry were present at all wavelengths and the biases did not cancel out as previously described, then our joint-fit would have given a time offset by this same value. If we then applied this offset time in our catwoman fits to the F322W2 spectroscopic data, we would not have observed any significant limb asymmetry. In this case, the assumed time of conjunction would be biased by an amount corresponding to the underlying asymmetry, which would have erased the asymmetry’s signal.
@@
To further evaluate our joint-fit’s effectiveness, we tested fitting each ancillary data set individually using a uniform-limb model and allowing to freely vary. We then compared the difference between each individually fit and the joint-fit . As just mentioned, when fitting the broadband JWST/NIRCam F322W2 data, we get a best-fit that 6.64 1.91 seconds later than the joint-fit result from the broadband light curve, and we also see chromatic offsets in the best-fit shown in Supplementary Figure 4 from the spectroscopic data. Both the SOAR and Spitzer/IRAC Ch.2 observations had differences consistent with zero (7.6 19.7 s and 4.42 7.39 s, respectively). The TESS data, which overlaps with the SOAR bandpass, had a best-fit offset of -33.5 19.0 s. This is negative and non-zero at over 1 significance, but it is not clear whether such large offset is real given the sparse 30-minute cadence of the data. The JWST/NIRCam F210M data, which was observed simultaneously with the NIRCam F322W2 data with the same cadence, had an offset of -1.11 1.06 s. This is also negative but nearly consistent with zero. Finally, the JWST/MIRI LRS data had a difference of 2.62 1.28 s which is non-zero by just over 2. Due to their superior data quality, our joint-fit is most effected by the JWST/NIRCam F210M and JWST/MIRI LRS observations. These individual fits show that both data sets may also be subject to some limb asymmetry, though not to the same degree as the JWST/NIRCam F322W2 data, and potentially of opposite polarity. Our multi-wavelength approach will therefore work as described to mitigate these individual biases. All together, these fits underscore the importance of combining data at multiple wavelengths. The mean offset of these individual fit values, weighted by their uncertainties (), is 0.41 s which is well within the joint-fit value’s uncertainty. Even if we ignore the largest offset, which was the TESS value, the mean offset is still only 0.47 s. To the best of our current knowledge, the we derive from our joint-fit is as accurate and bias-free as possible.
@@ 3.6 Forward Modeled Transmission Spectra
From the grid of models described above, we used an atmospheric metallicity of 10 times the solar value and carbon-to-oxygen ratio of 0.35 based on the best-fit retrieval values on HST/WFC3 and JWST/MIRI observations of this planet from ref.[26]. These values are also consistent with the additional retrievals in ref.[27]. We expect the metallicity and carbon-to-oxygen ratio to be globally uniform quantities, and do not vary them between the models for each limb. In this work, we do not seek to make specific, precise constraints on global properties, but rather develop an understanding for what may be causing the observed difference between the evening and morning limbs. To do this, we performed a grid fit over the heat redistribution parameter, over its full range, and the gray cloud opacity, ranging from 10-2.58 to 10-0.78 cm2 g-1. This heat redistribution parameter is effectively the temperature at the terminator, related by . We computed the value between the grid model and the observed morning and evening spectra individually, following the -minimization technique of refs.[42, 43].
@@ Appendix B Supplementary Information
Data Availability
The JWST/NIRCam (JWST GTO programme 1185; P.I. Greene; Obs. 8 and 9), JWST/MIRI (JWST GTO programme 1280; P.I. Lagage; Obs. 1), and TESS (P.I. Caldwell; Obs. I.D. hlsp_tess-spoc_tess_phot_0000000429302040-s0010_tess_v1) data are publicly available from the Mikulski Archive for Space Telescopes (MAST; https://mast.stsci.edu). The Spitzer/IRAC (Program 13052; P.I. Werner; AORKEY 62712320) data is publicly available from the Spitzer Heritage Archive (https://irsa.ipac.caltech.edu/applications/Spitzer/SHA/). The SOAR (Program N23A-840705; P.I. Murphy) data is available from Zenodo (https://zenodo.org/records/12747273).
Code Availability
The three reduction pipelines used in this work (Eureka!, t-shirt, and Pegasus) are all either currently open-source or are planned to be made open-source. The codes of each are hosted on Github, and links for each may be found in the Methods section.
Acknowledgements
Author Contributions Statement
M.M. led this analysis including formulating the idea, performing the data and supporting analyses, and was the primary author of the paper. T.G.B. contributed the Pegasus reduction of the NIRCam data, E.S. contributed the t-shirt reduction, and T.J.B. contributed the Eureka! reduction of the NIRCam F322W2 spectra and the broadband reduction of the MIRI/LRS data. T.J.B., E.S., and T.G.B. also contributed to the interpretation of our results and the writing of the paper. M.L. contributed the modeling analysis using ScCHIMERA. T.P.G. selected the object, devised the observing strategy, and contributed to the analysis. V.P., E.R., J.F., M.L., and L.W. each contributed to the methodology and interpretation of the results. M.R. designed the JWST/NIRCam instrument used for the observations, and provided the observing time necessary for completing these observations.
Competing Interests Statement
The authors declare no competing interests.